• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:09:34 on November 30, 2007

    Entry number 217266

    keyword=Shift Summary

    Day Shift 11/30/07 John LeRose SL Yi Zhang TO At start of shift runs 21300 & 3070 in progress from previous shift Stop at 08:17 and go to contolled access Left Q2 work, left arm polarity change, move spectrometers to 60°. Set right arm to 0.62267 GeV/c (quads cycled) About to come out of controlled access at end of shift.

    NOTE for future shifts: DO NOT cycle Q2 left! The power supply is one very sick puppy.